7. Dorman Industries has a new project available that requires an initial investment of $5.1 million. The project will provide unlevered cash flows of $735,000 per year for the next 20 years. The company will finance the project with a debt-to-value ratio of .4. The companys bonds have a YTM of 6.6 percent. The companies with operations comparable to this project have unlevered betas of 1.21, 1.14, 1.36, and 1.31. The risk-free rate is 3.6 percent, and the market risk premium is 6.8 percent. The company has a tax rate of 34 percent What is the NPV of this project? (Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.] NPV $ -701036 eBook &Resources eBook: 18.7 Beta and Leverage Check my work Previous attempt
Answer #1

First we need to compute the Levered beta of norman industries, in order to compute the cost of equity through the CAPM Model.

Levered Beta = Unlevered beta *( 1 + (1 - Tax rate) *D/E)

For the value of Unlevered beta, we take an arithmetic average of the unlevered beta of similar companies in the industry.

Which is approximately 1.255.

Putting it in the previous formula.

L Beta = 1.255 (1 + (1- TR) *D/E)

= 1.255 (1+ (1-0.34) *0.4)

= 1.58632

Ke = Rf + B (Rm - RF)

= 3.6 + 1.58632*(6.8)

= 14.386976%

KD = Cost of debt (1-tax rate)

= 6.6(1-.34) = 4.36%

Now, calculating the Weights of debt and equity from Debt to equity ratio.

D/E ratio = 0.4 or 4/10

That means, For every 10 equity there are 4 debt.

That means, Weight of Equity is 10/14 or 71% and 29%(4/14) for debt approximately.

So, the WACC = We * Ke + Wd * Kd

= 0.71* 0.14386976 + 0.29 * 0.0436

= 0.10214753 + 0.012644 = 11.479153 %

To calculate the NPV, Now we use the WACC we just calculated, and the NPV calculation are shown in the following screenshot.

Paste Blu Format Years Project X 20 $(51,00,000.00) 10 8 11 9 12 10 $ 13 11 14 12 15 13 $ 16 14 $ 17 15 $ 18 16 $ 19 17 20 18

Use the =NPV(rate, values) -initial cash flow forumula in excel to calculate this.